For this problem:
(-m^2n^-1)^-2/(mn)^-1
simplify
I get -n^3/m^3
but the books answer is
n^3/m^3
Am I not to assume a -1 in front of the first m in the numerator?
Please explain.

Answer by stanbon(75887)

(-m^2n^-1)^-2/(mn)^-1
simplify
--------------
(mn)/(-m^2n^-1)^2
--------
= (mn)/(m^4*n^-2)
--------
= (mn^3/m^4)
--------
= n^3/m^3
--------------
Cheers,
Stan H.
------------------